Grade 10Mechanics

A boat moves in a river of velocity 30m/s .and its motor propels it with a velocity of 45m/s in a direction perpendicular to the current. Find the resultant velocity of the boat

Jessica
6 Years agoGrade 10

1 Answer

Arun
6 Years ago
Dear studen
 
Resultant Velocity = R( let)
 
hence R^2 = 30^2 + 45^2 = 2925
 
R = sqrt(2925) = 54.08 m/s
 
Hope it helps